Contents

1 Here is a chapter title 3 1.1 Section header . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Another Section header . . . . . . . . . . . . . . . . . . . . . . . 4 1.3 An indexed list environment . . . . . . . . . . . . . . . . . . . . . 4 2 CONTENTS

chapter section subject notation sin@$\sin$

Chapter 1

Here is a chapter title

1.1 Section header

Here is some text.

Here is some more text.

4 CHAPTER 1. HERE IS A CHAPTER TITLE

more notation text section2 index second␣item

Here is some more text. Here is yet more text.

1.2 Another Section header

And here is some math: x_{b}^{1}.

Here is an index entry inside an fbox Here is an entry in a box.

1.3 An indexed list environment

1. First item

2. Second item

3. Third item 1.3. AN INDEXED LIST ENVIRONMENT 5

                                                                                 fifth␣item

4. Fourth item

5. Fifth item

6. Sixth item

Index

more, 4 subject, 3